OpenStudy (anonymous):

What is the axis of symmetry of the graph of y = –(x + 3)2 – 6?

OpenStudy (anonymous):

**y = –(x + 3)^2 – 6

OpenStudy (anonymous):

A. x = 6 B. x = –3 C. x = 1 D. x = –1

OpenStudy (anonymous):

y = -x^2 -6x + 9 - 6 y = -x^2 -6x + 3 x vertex : -b/2a 6 / -2 = -3 so x= -3 :)

OpenStudy (anonymous):

Thank you ((:

OpenStudy (phi):

y = –(x + 3)^2 – 6 @Coolsector: when the quadratic is in this form, you can read off the axis of symmetry. It's the number that makes (x+3) zero. i.e. x= -3. Also, quadratic peak (or minimum, depending on the sign) is (-3, -6) where you read off the y value -6.